write the equation using ax+by=c. the line passes throughthe points (3,5) and (-6,-7)
Answer by lwsshak3(11628)

write the equation using ax+by=c. the line passes through the points (3,5) and (-6,-7)
Start with form for equation of a line: y=mx+b, m=slope, b=y-intercept
slope, m=∆y/∆x=(5-(-7))/(3-(-6)=(5+7)/(3+6)=12/9=4/3
equation:
y=4x/3+b
solve for b using coordinates of one of the given points (3,5)
5=4*3/3+b
b=1
y=4x/3+1
3y=4x+3
4x-3y=-3